OpenStudy (anonymous):

Find y' for the following functions. You do not need to simplify your answers or show your work. (a) y(x) = integral (1 to x^2) (t^3 + t^2)dt (b) y(x) = exp(x) integral (0 to 1) cosh(x)dx

OpenStudy (anonymous):

Here's what it looks like in an equation: (a) \[y(x) = \int\limits_{1}^{x^2}(t^3+t^2)dt\] (b) \[y(x) = \exp(x) \int\limits_{0}^{1}\cosh(x)dx\]